[Brahms-l] Shift report 20040421 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Wed Apr 21 2004 - 07:57:08 EDT
------------------------------------------------------------
Date: Wed Apr 21 07:57:07 2004, Shift: 00:00-08:00
Supervisor: EBJ, with: 
Log book#15 pages 107 - 107

Shift Summary:
------------------------------------------------------------
1. RHIC delivered beam to Phenix and Star as soon as I arrived to the counting house at midnight.  Brahms didn't have collisions until 3:22am.

2.  I completed these runs at FS: 8 deg
 100k trig 6 evts for D1=D3=D4=0, D2= 1/4
 100k trig 6 evts for D2=D3=D4=0, D1= 1/4 

So far I have collected at FS: 4 deg.
  170k trig 6 evts for D1=D2=D3=D4=1/4 B, D5=350 (2) A 

3. See the Elog for TOFW and TOFW2 problems. I also did not see the bright blue light for the silicon preamp power when I did the check list.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Apr 21 07:58:16 2004

This archive was generated by hypermail 2.1.8 : Wed Apr 21 2004 - 07:58:38 EDT